OpenStudy (anonymous):

A farmer has a fence that encloses a square plot with an area of 100 square meters. If the farmer uses this fence to enclose a circular flower garden, what will the area of the garden be? (Use π = 3.14. Round your answer to two decimal places.)

OpenStudy (anonymous):

answer in meters squared

OpenStudy (anonymous):

p = 2*pi*r 100 meters divide by pi then divide by 2 is 50/pi meter (radius) A= Pi*r^2 (50/pi)^2 * pi = 2500/pi square meter in real number: 795.77 square meter

OpenStudy (anonymous):

answer was wrong

OpenStudy (anonymous):

so the length of the fence is 4*sqrt(100)=40

OpenStudy (anonymous):

*facepalm my fault completely. 100 square square garden is equal to 10 meter per side. then the parameter is 40 meter. using 40/2/3.14=6.394 = r 6.3694*3.14*6.3694 = 127.39 square meter.